The equation y= -16x+ 580 represents the population of mongoose in a particular area, where y is the mongoose population and x is the month. After 6 months, a population survey showed that there are 448 mongooses. What is the residual when compared to the model
Answer by jim_thompson5910(35256)

y = -16x+ 580


y = -16(6) + 580


y = -96 + 580


y = 484


The estimated population after 6 months is 484


The actual population measured after 6 months is 448


Residual: (observed) - (estimated) = (448) - (484) = -36